Shift report 20000717 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Jul 17 2000 - 23:55:57 EDT

  • Next message: http account: "Shift report 20000718 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Jul 17 23:55:57 2000, Shift: 16:00-24:00
    Supervisor: ykl, with: yb
    Log book#3 pages 24 - 25
    
    Shift Summary:
    ------------------------------------------------------------
    RHIC power supply studies continued until 21:40; beam
    was being retored after 22:30. Beam injection was
    very briefly successful and there is no beam at this
    moment.
    
    
    KO updated CAMAC readout to read new ZDC ADC data.
    
    Run 1805 is now in progress for a diagnostic pulser
    run for MM (ZDC) for 10000 events.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Jul 17 2000 - 23:59:07 EDT